OpenStudy (diyadiya):

Function \[f(x)= \frac{\sin^4x+\cos^4x}{x+x^tanx} \] is A.Even B.Odd C.Periodic with period pi D.Periodic with period 2pi

OpenStudy (anonymous):

rewrite the function

OpenStudy (diyadiya):

done

OpenStudy (experimentx):

seems periodic with 2pi .. but i wouldn't exactly call that

OpenStudy (diyadiya):

can you explain?

OpenStudy (experimentx):

definitely not even and odd, since f(-x) != -f(x) and f(-x) != f(x) either

OpenStudy (anonymous):

\[(\sin ^4 (-x) + \cos ^4 (-x)) /[ -x +(-x) \tan(-x)]\] numerator is equal to \[(sin4(x)+cos4(x))\] denominator, we know that tan (-x) = -tan(x) so -x + (-x) (- tan(x))= -x + tan(x) = so this will mean that the function is not odd or even

OpenStudy (zarkon):

is this the function \[f(x)= \frac{\sin^4x+\cos^4x}{x+x^{\tan(x)}}\]?

OpenStudy (experimentx):

the sin and cosines have period of 360, they are squares ... so 180 would be fine ... but there's tan involved ... so period would be no way 180

OpenStudy (zarkon):

If it is ... the answer is E: none of the above

OpenStudy (experimentx):

now if you plot the graph ... tan(x) would be inf at 90, so at odd multiple of 90's , graph would return to 0 .. at odd multiple of ... odd multiples of 90's

OpenStudy (experimentx):

oh man ... that would make period pi

OpenStudy (anonymous):

Graph the function ((sin(x))^4 + (cos(x))^4 )/ (x + x*tan(x)) in the http://www.coolmath.com/graphit/ I guess your function is wrong. check again

OpenStudy (experimentx):

looks like i have to check again the definition of periodicity .... http://en.wikipedia.org/wiki/Periodic_function can positive side of this graph be called periodic http://www4c.wolframalpha.com/Calculate/MSP/MSP94931a0iagba3h605ia400001b76666b87b9h4aa?MSPStoreType=image/gif&s=45&w=320&h=127&cdf=RangeControl

OpenStudy (experimentx):

if you can call it periodic then your answer would be C

OpenStudy (diyadiya):

Oh sorry i lost connection in between Function is \[\huge f(x)= \frac{\sin^4x+\cos^4x}{x+x^2tanx}\] I'm so sorry!

OpenStudy (diyadiya):

@experimentX @Zarkon @alireza.safdari

OpenStudy (experimentx):

i will give you the similar answer, if you can call it periodic, then the answer would be D .. ie period 2pi http://www.wolframalpha.com/input/?i=y+%3D+%28%28sinx%29%5E4%2B%28cos%29%5E4%29%2F%28x%2Bx%5E2*%28tanx%29%29

OpenStudy (diyadiya):

I'm confused :S

OpenStudy (anonymous):

It is an odd function continue from last post denominator, we know that tan (-x) = -tan(x) so -x + (-x) (- tan(x))= -x + ((-x)^2) *(-tan(x)) = -x - (x^2)(tanx) then we factor -1 (x + (x^2)(tanx)) now we know that f(x) = -f(-x)

OpenStudy (anonymous):

Diyadiya did you understand?

OpenStudy (experimentx):

Oo.. he's right

OpenStudy (diyadiya):

I'm going through it just a min

OpenStudy (experimentx):

odd function

OpenStudy (diyadiya):

Ok yeah i got it :) Thank you so much :)